# BSMamba: Brightness and Semantic Modeling for Long-Range Interaction in Low-Light Image Enhancement

### v1

@@ Abstract
Current low-light image enhancement (LLIE) methods face significant limitations in simultaneously improving brightness while preserving semantic consistency, fine details, and computational efficiency. With the emergence of state-space models, particularly Mamba, image restoration has achieved remarkable performance, yet existing visual Mamba approaches flatten 2D images into 1D token sequences using fixed scanning rules, critically limiting interactions between distant tokens with causal relationships and constraining their ability to capture meaningful long-range dependencies. To address these fundamental limitations, we propose BSMamba, a novel visual Mamba architecture comprising two specially designed components: Brightness Mamba and Semantic Mamba. The Brightness Mamba revolutionizes token interaction patterns by prioritizing connections between distant tokens with similar brightness levels, effectively addressing the challenge of brightness restoration in LLIE tasks through brightness-guided selective attention. Complementing this, the Semantic Mamba establishes priority interactions between tokens sharing similar semantic meanings, allowing the model to maintain contextual consistency by connecting semantically related regions across the image, thus preserving the hierarchical nature of image semantics during enhancement. By intelligently modeling tokens based on brightness and semantic similarity rather than arbitrary scanning patterns, BSMamba transcends the constraints of conventional token sequencing while adhering to the principles of causal modeling. Extensive experiments demonstrate that BSMamba achieves state-of-the-art performance in LLIE while preserving semantic consistency.
1 Introduction
@@
Images captured in low-light conditions often suffer from low contrast, color distortion, detail loss, and even complete information loss. LLIE aims to restore brightness and visual content in such images. Its advancements have significantly expanded its applications in fields like autonomous driving, security surveillance, and multimedia [33, 16, 7, 25].
@@
Traditional LLIE methods, including histogram equalization [23], gamma correction [19], and Retinex-based approaches [26, 47], typically struggle with extreme lighting conditions due to generalization issues, often neglecting noise and detail degradation, resulting in unsatisfactory performance. Deep learning-based LLIE methods [2, 45, 43, 56] attempt to learn the mapping between low-light and normal-light images through neural networks, demonstrating performance and visual quality superior to traditional approaches. However, most deep learning-based LLIE methods [35, 18, 50, 46] overlook brightness consistency, color fidelity, and fine image details during enhancement, leading to unpleasant color distortion, detail loss, and suboptimal brightness. Additionally, constrained by parameter efficiency, many LLIE methods [52, 45, 24, 43] cannot maintain both high performance and computational efficiency.
@@
Recently, state-space models [9] have developed rapidly as alternatives to CNNs and Transformers, offering significant advantages in long sequence modeling while reducing computational complexity from quadratic to linear. Visual Mamba-based methods [54, 21, 13, 41] have demonstrated superior performance over dominant Transformer models across various vision tasks, including image segmentation, super-resolution, and LLIE. However, applying Mamba to the visual domain still faces two critical limitations:
@@
First, while causal relationships between tokens in language naturally align with state-space model scanning mechanisms, in 2D images, tokens with causal relationships may be spatially distant, weakening the modeling capabilities between causally related tokens and limiting long-range modeling and generalization [44, 12]. Furthermore, flattening images into sequences disrupts the spatial dependencies and local adjacencies inherent in 2D images. This fundamental difference between language and vision domains creates challenges for adapting Mamba to visual tasks.
@@
Second, current visual Mamba-based LLIE methods [55, 1, 41] generally flatten 2D images into 1D sequences using fixed scanning rules, as illustrated in Fig. 1(a) for Wave-Mamba [55]. To compensate for interactions between adjacent and distant tokens, these methods often employ multi-directional forward and reverse scanning strategies. However, MambaIR [13, 12] has demonstrated that such multi-directional scanning contains significant information redundancy and fails to mitigate the decay of interactions between distant tokens. Various visual Mamba [20, 31] approaches have attempted to address this by dividing images into multiple windows or clusters (Fig. 1(b)), with each processed separately to ensure causally related tokens are handled together. Nevertheless, these approaches still lack flexibility in handling various object scales, and the problem of long-range decay between causally related tokens remains unresolved.
@@
To address these limitations, we introduce BSMamba, a novel Mamba architecture for LLIE tasks. BSMamba comprises Brightness Mamba and Semantic Mamba components. Brightness Mamba models the brightness essence of LLIE tasks by treating brightness levels as causal attributes between pixels, encouraging preferential interactions between spatially distant pixels with similar brightness properties, thereby enhancing long-range brightness-based causal interactions. Semantic Mamba leverages instance segmentation results to generate attention maps. Unlike conventional attention mechanisms that treat all regions equally, Semantic Mamba dynamically assigns attention scores based on segmentation confidence scores while maintaining the relative importance of features within each instance. Semantic Mamba unfolds the image into a 1D sequence according to each object’s attention score, encouraging priority interactions between distant tokens with similar semantic scores on a semantically consistent basis. Brightness Mamba and Semantic Mamba in BSMamba require only two scanning passes, improving efficiency while effectively enhancing brightness and semantic interactions between distant tokens, as shown in Fig. 1(c). After processing by BSMamba, the predicted image is finally fed into a Detail Enhancement Network (DE-Net) to further optimize image edge details. The overall framework follows a three-step strategy: ensuring brightness, semantic, and detail consistency throughout the LLIE process.
@@
2 Related Work
2.1 Low-Light Scene Enhancement
@@
Traditional LLIE methods, such as histogram equalization [23], gamma correction [19], and Retinex theory [26], often introduce artifacts and noise under extreme or complex lighting conditions. With the rise of deep learning, neural network-based LLIE methods [45, 48, 46] have become dominant. Retinex-based approaches [8, 42], like Retinexformer [2], combine Retinex theory with Transformers to address long-range dependencies, while IAGC [38] introduces learnable gamma correction for adaptive enhancement. In addition, many methods incorporate priors to guide LLIE. For instance, SMG-LLE [46] uses structural maps for enhancement, SKF [43] integrates semantic priors, and LEDN [40] incorporates depth information for improved restoration. However, these methods face challenges such as errors from pre-trained models and increased computational costs due to prior generation modules. Recently, frequency-domain LLIE methods have emerged. FourLLIE [35] enhances brightness via amplitude mapping in Fourier space, while DMFourLLIE [50] jointly enhances phase and amplitude using infrared modalities and brightness attention. WaveMamba [55] applies state-space models in the wavelet domain for brightness enhancement and noise removal, and WalMaFa [34] combines wavelet-based Mamba with Fourier adjustments. However, these methods rely on vanilla Mamba and fail to address its limitations in modeling long-range interactions in 2D images.
2.2 State-Space Models
@@
State-space models (SSMs) have gained attention for addressing long-range dependencies, as demonstrated by LSSL [11] and S4 [10], which serve as alternatives to CNNs and Transformers. Mamba [9], a selective SSM, has been extended to visual tasks like classification [54], restoration [13, 5], and fusion [27]. In LLIE, RetinexMamba [1] combines Retinexformer with SSMs for faster processing, while WaveMamba [55] and WalMaFa [34] leverage Mamba for low-frequency feature extraction and global brightness modeling. However, these methods rely on vanilla Mamba, which struggles with weakened long-range interactions in 2D images. To address this, improved methods have been proposed. LocalMamba [20] uses window-based scanning for better local detail representation, QuadMamba [44] employs quadtree-based scanning to preserve spatial locality, and ARM clusters adjacent blocks for efficient scanning. LLEMamba [51] adopts bidirectional scanning to balance local and global focus. However, these approaches fail to model causal relationships between pixels, limiting effective interactions and overall performance.
3 Method
3.1 Overview
@@
The proposed BSMamba framework is designed to address the challenges of LLIE by simultaneously modeling brightness and semantic consistency. As illustrated in Fig. 2, the input low-light image is first processed through an initial convolutional layer to extract features. These feature maps are then refined by a series of BSMamba blocks, where the Brightness Mamba and Semantic Mamba modules enhance the interactions between causally related tokens based on brightness and semantic hierarchies. The refined feature maps are subsequently passed through a Detail Enhancement Network (DE-Net) to reconstruct spatial details, producing the final enhanced image. Below, we provide a detailed description of each component.
@@
3.2 Brightness Mamba
To overcome the fundamental limitations of existing visual Mamba methods in modeling long-range dependencies, we propose Brightness Mamba [9, 55], a novel component that prioritizes interactions between distant tokens with similar brightness levels. Specifically, Brightness Mamba utilizes DA-CLIP [30] to compute brightness grading scores by calculating similarity scores between all image pixels and brightness descriptors, reflecting the brightness levels of all pixels across the image. We unfold the image into a 1D sequence based on brightness scores, where pixels with similar brightness levels are grouped together, promoting effective interaction between spatially distant pixels with similar brightness characteristics.
@@ 3.2 Brightness Mamba
where $\mathbf{X}$ is the input image. The BHS mechanism, as shown in Fig. 3, consists of two key steps: 1) brightness hierarchy computation using DA-CLIP [30], and 2) brightness hierarchy sorting.
Figure 4: Semantic Hierarchy Scanning (SHS) in Semantic Mamba. The SHS structure facilitates semantic-aware token interactions by grouping tokens based on their semantic significance, ensuring consistent enhancement within object instances.
@@
[…]
@@
where $\hat{\mathbf{X}}_{\text{sort}}$ represents the 1D sequence of tokens reordered based on their brightness scores, ensuring that pixels with similar brightness characteristics are processed sequentially, regardless of their spatial positions in the original image. $\hat{\mathbf{I}}_{\text{sort}}$ is the inverse index used to reorder the tokens back to their original positions. As shown in Fig. 3, tokens from different spatial locations but with similar brightness levels (indicated by blocks of similar colors) are grouped together in the sorted sequence. By prioritizing interactions between tokens with similar brightness levels, Brightness Mamba effectively models the causal relationships of brightness across the image space, addressing the core challenges of LLIE.
3.3 Semantic Mamba
While addressing brightness consistency is crucial for LLIE tasks, maintaining semantic coherence between object instances is equally essential to ensure perceptually natural results. To accomplish this, we propose the Semantic Mamba module, as shown in Fig.2, which leverages instance segmentation information to guide token interactions based on semantic relationships.
Similar to Brightness Mamba, Semantic Mamba incorporates the SHS mechanism within a state-space block structure. As shown in Fig. 4, the Semantic Hierarchy Scanning (SHS) mechanism consists of two key steps: (1) semantic mapping using Mask R-CNN [17], and (2) semantic hierarchy sorting.
@@ 3.3 Semantic Mamba
where $\mathbf{X}$ is the input image, ${G}_{\text{Mask-RCNN}}$ generates instance segmentation masks with confidence scores, and $\mathcal{M}_{\text{semantic}}$ maps these instances to a continuous semantic attention map $\mathbf{M}_{s}\in\mathbb{R}^{B\times H\times W}$.
The semantic mapping function $\mathcal{M}_{\text{semantic}}$ implements a dynamic grading strategy:
@@
Methods LSRW-Nikon #Param #FLOPs
PSNR ↑ SSIM ↑ LPIPS ↓ (M) (G)
MIRNet [48] 17.10 0.5022 0.2170 31.79 785.1
Kind++ [53] 14.79 0.4749 0.2111 8.27 -
SGM [47] 15.73 0.4971 0.2234 2.31 -
FECNet [18] 17.06 0.4999 0.2192 0.15 5.82
HDMNet [28] 16.65 0.4870 0.2157 2.32 -
FourLLIE [35] 17.82 0.5036 0.2150 0.12 4.07
UHDFour [24] 17.94 0.5195 0.2546 17.54 4.78
Retinexformer [2] 17.64 0.5082 0.2784 1.61 15.57
WaveMamba [55] 17.34 0.5192 0.2294 1.26 7.22
DMFourLLIE [50] 17.04 0.5294 0.2178 0.75 5.81
RetinexMamba [1] 17.59 0.5133 0.2534 3.59 34.76
Ours 18.04 0.5327 0.2012 0.68 8.15
Table 1: Quantitative comparison on LSRW-Nikon dataset. The best results are highlighted in bold.
@@
Methods LSRW-Huawei #Param #FLOPs
PSNR ↑ SSIM ↑ LPIPS ↓ (M) (G)
NPE [37] 17.08 0.3905 0.2303 - -
LIME [14] 17.00 0.3816 0.2069 - -
SRIE [8] 13.42 0.4282 0.2166 - -
Kind [52] 16.58 0.5690 0.2259 8.02 34.99
MIRNet [48] 19.98 0.6085 0.2154 31.79 785.1
Kind++ [53] 15.43 0.5695 0.2366 8.27 -
SGM [47] 18.85 0.5991 0.2492 2.31 -
FECNet [18] 21.09 0.6119 0.2341 0.15 5.82
HDMNet [28] 20.81 0.6071 0.2375 2.32 -
SNR-Aware [45] 20.67 0.5911 0.1923 39.12 26.35
FourLLIE [35] 21.11 0.6256 0.1825 0.12 4.07
UHDFour [24] 19.39 0.6006 0.2466 17.54 4.78
Retinexformer [2] 21.23 0.6309 0.1699 1.61 15.57
WaveMamba [55] 21.19 0.6391 0.1818 1.26 7.22
UHDFormer [36] 20.64 0.6244 0.1812 0.34 3.24
DMFourLLIE [50] 21.47 0.6331 0.1781 0.75 5.81
RetinexMamba [1] 20.88 0.6298 0.1689 3.59 34.76
Ours 21.49 0.6435 0.1613 0.68 8.15
Table 2: Quantitative comparison on LSRW-Huawei dataset. The best results are highlighted in bold.
@@
FourLLIE [35] 20.99 0.8071 0.0952 23.45 0.8450 0.0613
UHDFour [24] 22.89 0.8147 0.0934 27.27 0.8579 0.0617
SNR-SKF [43] 20.51 0.7110 0.1810 21.82 0.7468 0.1513
Retinexformer [2] 22.71 0.8177 0.0922 24.55 0.8434 0.0627
RetinexMamba [1] 23.15 0.8210 0.0876 27.31 0.8667 0.0551
WaveMamba [55] 22.76 0.8419 0.0791 27.87 0.8935 0.0451
DMFourLLIE [50] 21.58 0.8349 0.0821 26.74 0.8792 0.0528
@@
Table 3: Quantitative comparison on LOL dataset. The best results are highlighted in bold.
@@
The Semantic Mamba module complements the Brightness Mamba module within the BSMamba framework. While Brightness Mamba focuses on enhancing the brightness consistency of the image based on illumination-related features, Semantic Mamba ensures that the semantic relationships between different image regions are preserved during enhancement. This dual-perspective approach addresses both critical aspects of LLIE: Brightness Mamba establishes long-range dependencies between pixels with similar illumination conditions. Semantic Mamba maintains semantic coherence by ensuring consistent enhancement within object instances.
3.4 Detail Enhancement Network (DE-Net)
After processing with BSMamba blocks, which restore brightness relationships and semantic consistency, we further refine the spatial details and texture information through our proposed Detail Enhancement Network (DE-Net). As illustrated in Fig. 2, DE-Net is specifically designed to enhance fine-grained details that may remain partially degraded after the initial enhancement. The DE-Net adopts an encoder-decoder architecture with Fast Fourier Convolution (FFC) [3] blocks in the bottleneck. The FFC architecture divides feature channels into spatial and spectral branches, allowing the network to simultaneously capture local spatial details and global frequency patterns, which is particularly beneficial for enhancing texture and structural details [6].
@@ 3.4 Detail Enhancement Network (DE-Net)
where $\mathcal{E}$ represents the Canny edge extraction operation, $\mathbf{Y}$ is the predicted enhanced image, $\mathbf{Y}_{\text{gt}}$ is the ground truth image, and BCE denotes the binary cross-entropy loss function [32].
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{1}}+\lambda_{2}\mathcal{L}_{\text{ssim}},+\lambda_{3}\mathcal{L}_{\text{canny}}$ (10)
where loss weights $\lambda_{1},\lambda_{2},\lambda_{3}=[1.0,0.5,0.1]$, $\mathcal{L}_{\text{1}}$ is the $l_{1}$ loss, $\mathcal{L}_{\text{ssim}}$ denotes the structural similarity (SSIM) loss [39].
4 Experiments
Figure 8: Visual comparison on DICM (first and second rows) and NPE (third row) datasets. Zoomed-in provide a clearer comparison.

### v2

@@ Abstract
Current low-light image enhancement (LLIE) methods face significant limitations in simultaneously improving brightness while preserving semantic consistency, fine details, and computational efficiency. With the emergence of state-space models, particularly Mamba, image restoration has achieved remarkable performance, yet existing visual Mamba approaches flatten 2D images into 1D token sequences using fixed scanning rules, critically limiting interactions between distant tokens with causal relationships and constraining their ability to capture meaningful long-range dependencies. To address these fundamental limitations, we propose BSMamba, a novel visual Mamba architecture comprising two specially designed components: Brightness Mamba and Semantic Mamba. The Brightness Mamba revolutionizes token interaction patterns by prioritizing connections between distant tokens with similar brightness levels, effectively addressing the challenge of brightness restoration in LLIE tasks through brightness-guided selective attention. Complementing this, the Semantic Mamba establishes priority interactions between tokens sharing similar semantic meanings, allowing the model to maintain contextual consistency by connecting semantically related regions across the image, thus preserving the hierarchical nature of image semantics during enhancement. By intelligently modeling tokens based on brightness and semantic similarity rather than arbitrary scanning patterns, BSMamba transcends the constraints of conventional token sequencing while adhering to the principles of causal modeling. Extensive experiments demonstrate that BSMamba achieves state-of-the-art performance in LLIE while preserving semantic consistency. Code is available at https://github.com/bywlzts/BSMamba.
Introduction
@@
Images captured in low-light conditions often suffer from low contrast, color distortion, detail loss, and complete information loss. Low Light Image Enhancement (LLIE) aims to restore brightness and visual content in such images, significantly expanding its applications in autonomous driving, security surveillance, and multimedia (Hashmi et al. 2023; Du, Shi, and Deng 2024; Li et al. 2024a).
@@
Traditional LLIE methods, such as histogram equalization (Lee, Lee, and Kim 2013), gamma correction (Huang, Cheng, and Chiu 2012), and Retinex-based approaches (Li et al. 2018; Yang et al. 2021), struggle in extreme lighting conditions due to generalization issues, often neglecting noise and detail degradation, which leads to unsatisfactory performance. In contrast, learning-based LLIE methods (Cai et al. 2023; Xu et al. 2022; Zou et al. 2024b) learn the mapping between low-light and normal-light images through neural networks, offering superior performance and visual quality. However, many deep learning LLIE methods (Wang, Wu, and Zhi 2023; Huang et al. 2022; Zhang et al. 2024a; Xu, Wang, and Lu 2023) overlook brightness consistency, color fidelity, and fine details during enhancement, resulting in color distortion, detail loss, and suboptimal brightness. Furthermore, due to parameter constraints, numerous LLIE methods (Li et al. 2023; Wu et al. 2023) struggle to balance high performance with computational efficiency.
@@
Recently, state-space models (Gu and Dao 2023) have developed rapidly as alternatives to CNNs and Transformers, offering significant advantages in long sequence modeling while reducing computational complexity from quadratic to linear. Visual Mamba-based methods (Huang et al. 2024b; Guo et al. 2025; Weng et al. 2025) have demonstrated superior performance over dominant Transformer models across various vision tasks, including image segmentation, super-resolution, and LLIE. However, applying Mamba to the visual domain still faces two critical limitations:
@@
First, while causal relationships between tokens in language align well with state-space model scanning mechanisms, tokens with causal relationships in 2D images may be spatially distant, which weakens the modeling capabilities between these tokens and limits long-range modeling and generalization (Xie et al. 2025). Additionally, flattening images into sequences disrupts the spatial dependencies and local adjacencies intrinsic to 2D images. This fundamental difference between language and vision domains poses challenges for adapting Mamba to visual tasks.
@@
Second, current Mamba-based LLIE methods (Zou et al. 2024a; Bai, Yin, and He 2024; Weng et al. 2025) generally flatten 2D images into 1D sequences using fixed scanning rules, as illustrated in Fig. 1(a). To compensate for interactions between adjacent and distant tokens, these methods often employ multi-directional forward and reverse scanning strategies. However, MambaIR(Guo et al. 2025, 2024) has revealed that this approach introduces significant information redundancy and fails to effectively address the decay of interactions among distant tokens. Various visual Mamba techniques (Huang et al. 2024a; Ren et al. 2024) have attempted to tackle this issue by segmenting images into multiple windows or clusters (Fig. 1(b)), processing each separately to group causally related tokens. Despite these efforts, these methods still lack flexibility in managing diverse object scales, leaving the challenge of long-range decay between causally related tokens unresolved.
@@
To address these limitations, we introduce BSMamba, a novel Mamba architecture for LLIE tasks. BSMamba comprises Brightness Mamba and Semantic Mamba components. Brightness Mamba models the brightness essence of LLIE tasks by treating brightness levels as causal attributes between pixels, encouraging preferential interactions between spatially distant pixels with similar brightness properties, thereby enhancing long-range brightness-based causal interactions. Semantic Mamba leverages instance segmentation results to generate attention maps. Unlike conventional attention mechanisms that treat all regions equally, Semantic Mamba dynamically assigns attention scores based on segmentation confidence scores while maintaining the relative importance of features within each instance. Semantic Mamba unfolds the image into a 1D sequence according to each object’s attention score, encouraging priority interactions between distant tokens with similar semantic scores on a semantically consistent basis. Brightness Mamba and Semantic Mamba in BSMamba require only two scanning passes, improving efficiency while effectively enhancing brightness and semantic interactions between distant tokens, as shown in Fig. 1(c). After processing by BSMamba, the predicted image is finally fed into a Detail Enhancement Network (DE-Net) to further optimize image edge details.
@@
Related Work
Low-Light Scene Enhancement
@@
Traditional LLIE methods, such as histogram equalization, gamma correction, and Retinex theory, often introduce artifacts and noise under extreme lighting conditions. With the advent of deep learning, neural network-based LLIE methods (Zamir et al. 2020; Zhang et al. 2025; Yan et al. 2025) have become prevalent. Retinex-based approaches (Wu et al. 2022), including Retinexformer (Cai et al. 2023), combine Retinex theory with Transformers to better manage long-range dependencies, while IAGC (Wang et al. 2023) introduces learnable gamma correction for adaptive enhancement. Many methods also integrate priors to guide LLIE. For instance, SMG-LLE (Xu, Wang, and Lu 2023) uses structural maps, SKF (Wu et al. 2023) incorporates semantic priors, and LEDN (Wang et al. 2024b) adds depth information to enhance restoration. However, these approaches face challenges related to errors from pre-trained models and increased computational costs due to the prior generation. Recently, frequency-domain LLIE methods have emerged, such as FourLLIE (Wang, Wu, and Zhi 2023), which enhances brightness through amplitude mapping in Fourier space, and DMFourLLIE (Zhang et al. 2024a), which improves both phase and amplitude using infrared modalities and brightness attention. In addition, wave-based methods (Zou et al. 2024a; Tan et al. 2024) apply state-space models in the wavelet domain for brightness enhancement and noise removal. However, these methods rely on standard Mamba, which fails to effectively model long-range interactions in 2D images.
State-Space Models
@@
State-space models (SSMs) have gained traction for addressing long-range dependencies, as shown by LSSL (Gu et al. 2021) and S4 (Gu, Goel, and Ré 2021), which serve as alternatives to CNNs and Transformers. Mamba (Gu and Dao 2023), a selective SSM, has been adapted for visual tasks such as classification (Zhu et al. 2024), restoration (Guo et al. 2025; Dong et al. 2024), and fusion (Li et al. 2024b). In LLIE, RetinexMamba (Bai, Yin, and He 2024) merges Retinexformer with SSMs for improved processing speed, while WaveMamba (Zou et al. 2024a) and WalMaFa (Tan et al. 2024) utilize Mamba for low-frequency feature extraction and global brightness modeling. Yet, these methods still depend on standard Mamba, which struggles with long-range interactions in 2D images. To improve performance, several enhanced methods have been proposed. LocalMamba (Huang et al. 2024a) uses window-based scanning for better local detail, QuadMamba (Xie et al. 2025) employs quadtree-based scanning to maintain spatial locality, and ARM clusters adjacent blocks for efficient processing. LLEMamba (Zhang et al. 2024b) adopts bidirectional scanning to strike a balance between local and global focus. However, these approaches still struggle to model causal relationships between pixels, limiting effective interactions and overall performance.
Method
Overview
@@
The proposed BSMamba framework is designed to address the challenges of LLIE by simultaneously modeling brightness and semantic consistency. As illustrated in Fig. 2, the input low-light image is first processed through an initial convolutional layer to extract features. These feature maps are then refined by a series of BSMamba blocks, where the Brightness Mamba and Semantic Mamba modules enhance the interactions between causally related tokens based on brightness and semantic hierarchies. The refined feature maps are subsequently passed through a Detail Enhancement Network (DE-Net) to reconstruct spatial details, producing the final enhanced image.
@@
Brightness Mamba
To overcome the fundamental limitations of existing visual Mamba methods in modeling long-range dependencies, we propose Brightness Mamba, a novel component that prioritizes interactions between distant tokens with similar brightness levels. Specifically, Brightness Mamba utilizes DA-CLIP (Luo et al. 2023) to compute brightness grading scores by calculating similarity scores between all image pixels and brightness descriptors, reflecting the brightness levels of all pixels across the image. We unfold the image into a 1D sequence based on brightness scores, where pixels with similar brightness levels are grouped together, promoting effective interaction between spatially distant pixels with similar brightness characteristics.
@@
where $\mathbf{X}$ is the input image. The BHS mechanism, as shown in Fig. 3, consists of two key steps: 1) brightness hierarchy computation, and 2) brightness hierarchy sorting.
Figure 4: Semantic Hierarchy Scanning (SHS) in Semantic Mamba. SHS facilitates semantic-aware token interactions by grouping tokens based on their semantic significance, ensuring consistent enhancement within object instances.
@@
$\mathbf{M}_{b}=\text{Sim}(\Phi_{image}(\mathbf{X}),\Phi_{text}(\text{`low-light'})),$ (3)
@@
where $\hat{\mathbf{X}}_{\text{sort}}$ represents the 1D sequence of tokens reordered based on their brightness scores, ensuring that pixels with similar brightness characteristics are processed sequentially, regardless of their spatial positions in the original image. $\hat{\mathbf{I}}_{\text{sort}}$ is the inverse index used to reorder the tokens back to their original positions. Tokens from different spatial locations but with similar brightness levels (indicated by blocks of similar colors) are grouped together in the sorted sequence. By prioritizing interactions between tokens with similar brightness levels, Brightness Mamba effectively models the causal relationships of brightness across the image space, addressing the core challenges of LLIE.
Semantic Mamba
While addressing brightness consistency is crucial for LLIE tasks, maintaining semantic coherence between object instances is equally essential to ensure perceptually natural results. To accomplish this, we propose the Semantic Mamba module, as shown in Fig.2, which leverages instance segmentation information to guide token interactions based on semantic relationships. Similar to Brightness Mamba, Semantic Mamba incorporates the SHS mechanism within a state-space block structure. As shown in Fig. 4, the Semantic Hierarchy Scanning (SHS) mechanism consists of two key steps: 1) semantic mapping using Mask R-CNN (He et al. 2017), and 2) semantic hierarchy sorting.
@@
where $\mathbf{X}$ is the input image, ${G}_{\text{Mask-RCNN}}$ generates instance segmentation masks with confidence scores, and $\mathcal{M}_{\text{semantic}}$ maps these instances to a continuous semantic attention map $\mathbf{M}_{s}\in\mathbb{R}^{B\times H\times W}$. The semantic mapping function $\mathcal{M}_{\text{semantic}}$ implements a dynamic grading strategy:
@@
Methods Venue LSRW-Nikon #Param
PSNR ↑ SSIM ↑ LPIPS ↓ (M)
MIRNet ECCV, 20 17.10 0.5022 0.2170 31.79
Kind++ IJCV, 21 14.79 0.4749 0.2111 8.27
SGM TIP, 21 15.73 0.4971 0.2234 2.31
FECNet ECCV, 22 17.06 0.4999 0.2192 0.15
HDMNet MM, 22 16.65 0.4870 0.2157 2.32
FourLLIE MM, 23 17.82 0.5036 0.2150 0.12
UHDFour ICLR, 23 17.94 0.5195 0.2546 17.54
Retinexformer ICCV, 23 17.64 0.5082 0.2784 1.61
WaveMamba MM, 24 17.34 0.5192 0.2294 1.26
UHDFormer AAAI, 24 17.01 0.5186 0.2793 0.34
DMFourLLIE MM, 24 17.04 0.5294 0.2178 0.75
RetinexMamba ICONIP, 24 17.59 0.5133 0.2534 3.59
MambaLLIE NIPS, 24 17.25 0.5084 0.2612 2.28
CWNet ICCV, 25 17.38 0.5119 0.2732 1.23
CIDNet CVPR, 25 17.16 0.4975 0.2859 1.88
Ours 18.04 0.5327 0.2012 0.68
Table 1: Quantitative comparison on LSRW-Nikon dataset. The best and second-best results are highlighted in bold and underlined, respectively.
@@
Methods Venue LSRW-Huawei #FLOPs
PSNR ↑ SSIM ↑ LPIPS ↓ (G)
MIRNet ECCV, 20 19.98 0.6085 0.2154 785.1
Kind++ IJCV, 21 15.43 0.5695 0.2366 -
SGM TIP, 21 18.85 0.5991 0.2492 -
FECNet ECCV, 22 21.09 0.6119 0.2341 5.82
HDMNet MM, 22 20.81 0.6071 0.2375 -
SNR-Aware ICIP, 22 20.67 0.5911 0.1923 26.35
FourLLIE MM, 22 21.11 0.6256 0.1825 4.07
UHDFour ICLR, 23 19.39 0.6006 0.2466 4.78
Retinexformer ICCV, 23 21.23 0.6309 0.1699 15.57
WaveMamba MM, 24 21.19 0.6391 0.1818 7.22
UHDFormer AAAI, 24 20.64 0.6244 0.1812 3.24
DMFourLLIE MM, 24 21.47 0.6331 0.1781 5.81
RetinexMamba ICONIP, 24 20.88 0.6298 0.1689 34.76
MambaLLIE NIPS, 24 20.98 0.6388 0.1894 20.85
CWNet ICCV, 25 21.50 0.6397 0.1682 11.3
CIDNet CVPR, 25 20.30 0.6054 0.1973 7.57
Ours 21.49 0.6435 0.1613 8.15
Table 2: Quantitative comparison on LSRW-Huawei dataset.
@@
FourLLIE 20.99 0.8071 0.0952 23.45 0.8450 0.0613
UHDFour 22.89 0.8147 0.0934 27.27 0.8579 0.0617
SNR-SKF 20.51 0.7110 0.1810 21.82 0.7468 0.1513
Retinexformer 22.71 0.8177 0.0922 24.55 0.8434 0.0627
RetinexMamba 23.15 0.8210 0.0876 27.31 0.8667 0.0551
WaveMamba 22.76 0.8419 0.0791 27.87 0.8935 0.0451
DMFourLLIE 21.58 0.8349 0.0821 26.74 0.8792 0.0528
MambaLLIE 22.80 0.8315 0.8450 26.96 0.8875 0.0483
CIDNet 23.81 0.8574 0.0856 - - -
CWNet 23.60 0.8496 0.0648 27.39 0.9005 0.0383
@@
Table 3: Quantitative comparison on LOL dataset.
@@
The Semantic Mamba complements the Brightness Mamba within the BSMamba framework. While Brightness Mamba focuses on enhancing the brightness consistency of the image based on illumination-related features, Semantic Mamba ensures that the semantic relationships between different image regions are preserved during enhancement. This dual-perspective approach addresses both critical aspects of LLIE: Brightness Mamba establishes long-range dependencies between pixels with similar illumination conditions. Semantic Mamba maintains semantic coherence by ensuring consistent enhancement within object instances.
Detail Enhancement Network (DE-Net)
After processing with BSMamba blocks, which restore brightness relationships and semantic consistency, we further refine the spatial details and texture information through our proposed Detail Enhancement Network (DE-Net). As illustrated in Fig. 2, DE-Net is specifically designed to enhance fine-grained details that may remain partially degraded after the initial enhancement. The DE-Net adopts an encoder-decoder architecture with Fast Fourier Convolution (FFC) (Chi, Jiang, and Mu 2020) blocks in the bottleneck. The FFC architecture divides feature channels into spatial and spectral branches, allowing the network to simultaneously capture local spatial details and global frequency patterns, which is particularly beneficial for enhancing texture and structural details (Dong, Cao, and Fu 2022).
@@
where $\mathcal{E}$ represents the Canny edge extraction operation, $\mathbf{Y}$ is the predicted enhanced image, $\mathbf{Y}_{\text{gt}}$ is the ground truth image, and BCE denotes the binary cross-entropy loss function (Ruby, Yendapalli et al. 2020).
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{1}}+\lambda_{2}\mathcal{L}_{\text{ssim}}+\lambda_{3}\mathcal{L}_{\text{canny}},$ (10)
where loss weights $\lambda_{1},\lambda_{2},\lambda_{3}=[1.0,0.5,0.1]$, $\mathcal{L}_{\text{1}}$ is the $l_{1}$ loss, $\mathcal{L}_{\text{ssim}}$ denotes the SSIM loss (Wang et al. 2004).
Experiments